A viscous fluid of density 1050 kg m−3 and viscosity 12×10−4 Pa s flows through a pipe with a diameter of 5.0 cm.
What is the maximum speed at which the fluid can flow where the flow can be guaranteed to be laminar? (to 2 s.f and in m s−1)
1
Expert's answer
2020-04-27T10:21:02-0400
"Re=\frac{\rho dv}{\eta}=10"

"\frac{(1050)(0.05)v}{12\cdot10^{-4}}=10"

"v=2.3\cdot10^{-4}\frac{m}{s}"